# Composing Global Optimizers to Reasoning Tasks via Algebraic Objects in Neural Nets

### v2

@@ Abstract
We prove rich algebraic structures of the solution space for 2-layer neural networks with quadratic activation and $L_{2}$ loss, trained on reasoning tasks in Abelian group (e.g., modular addition). Such a rich structure enables analytical construction of global optimal solutions from partial solutions that only satisfy part of the loss, despite its high nonlinearity. We coin the framework as CoGO (Composing Global Optimizers). Specifically, we show that the weight space over different numbers of hidden nodes of the 2-layer network is equipped with a semi-ring algebraic structure, and the loss function to be optimized consists of monomial potentials, which are ring homomorphism, allowing partial solutions to be composed into global ones by ring addition and multiplication. Our experiments show that around $95\%$ of the solutions obtained by gradient descent match exactly our theoretical constructions. Although the global optimizers constructed only required a small number of hidden nodes, our analysis on gradient dynamics shows that overparameterization asymptotically decouples training dynamics and is beneficial. We further show that training dynamics favors simpler solutions under weight decay, and thus high-order global optimizers such as perfect memorization are unfavorable. The code is at [^1] https://github.com/facebookresearch/luckmatters/tree/yuandong3/ssl/real-dataset.
@@ 1 Introduction
As a result, our theoretical framework, named CoGO (i.e., Composing Global Optimizers), successfully constructs two distinct types of Fourier-based global optimizers of per-frequency order 4 (or “$2\times 2$”) and order 6 (or “$2\times 3$”), and a global optimizer of order $d^{2}$ that correspond to perfect memorization. Empirically, we demonstrate that around $95\%$ of the solutions obtained from gradient descent (with weight decay) have the predicted structure and match exactly with our theoretical construction of order-4 and order-6 solutions. In addition, we also analyze the training dynamics, and show that the dynamics favors low-order global optimizers, since global optimizers algebraically connected by ring multiplication can be proven to also be topologically connected. Therefore, high-order solution like perfect memorization is unfavorable in the dynamics. When the network width goes to infinite, the dynamics of monomial potentials becomes decoupled, demystifying why overparameterization improves the performance.
@@ 2 Related Works
Figure 1: Overview of proposed theoretical framework CoGO. (1) The family of 2-layer neural networks, $\mathcal{Z}$, form a semi-ring algebraic structure (Theorem 2) with ring addition and multiplication (Def. 5). $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$ where $\mathcal{Z}_{q}$ is a collection of all weights with order-$q$ (i.e., $q$ hidden nodes). (2) For outcome prediction of Abelian group multiplication, the MSE loss $\ell({\bm{z}})$ is a function of monomial potentials (MPs) $r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$ (Theorem 1), which are ring homomorphism (Theorem 3). (3) Thanks to the property of ring homomorphism, global optimizers to MSE loss $\ell({\bm{z}})$ with quadratic activation can be constructed algebraically from partial solutions that only satisfy a subset of constraints (Sec. 5.1) using ring addition and multiplication, instead of running gradient descent. Examples include Fourier solution ${\bm{z}}_{F6}$ (Corollary 2) and ${\bm{z}}_{F4/6}$ (Corollary 4) and perfect memorization solution ${\bm{z}}_{M}$ (Corollary 3). In Sec. 6, we analyze the role played of MPs in gradient dynamics, showing that the dynamics favors low-order global optimizers (Theorem 6) under weight decay regularization, and the dynamics of MPs become decoupled with infinite width (Theorem 7). In Sec. 7 we show that the gradient descent solutions match exactly with our theoretical construction.
@@ 3 Decoupling $L_{2}$ Loss for reasoning tasks of Abelian group
Basic Ring theory. A set $\mathcal{Z}$ forms a ring, if there exists two operations, addition $+$ and multiplication *, so that (1) $\langle\mathcal{Z},+\rangle$ forms an Abelian group, (2) $\langle\mathcal{Z},*\rangle$ is a monoid (i.e., a group without inverse), and (3) multiplication distributes with addition (i.e., $a*(b+c)=a*b+a*c$ and $(b+c)*a=b*a+c*a$). $\mathcal{Z}$ is called semi-ring if $\langle\mathcal{Z},+\rangle$ is a monoid.
@@
$\ell=\sum_{i}\Big{\|}P^{\perp}_{1}\left(\frac{1}{2d}{\bm{o}}[i]-l[i]\right)\Big{\|}^{2},\quad\quad{\bm{o}}[i]=V\sigma(W^{\top}{\bm{f}}[i])=\sum_{j}{\bm{v}}_{j}\sigma(\mathbf{w}_{j}^{\top}{\bm{f}}[i])$ (1)
where $\sigma(x)=x^{2}$ is the quadratic activation function (Du & Lee, 2018; Allen-Zhu & Li, 2023), $P_{1}^{\perp}=I-\frac{1}{d}{\bm{1}}{\bm{1}}^{\top}$ is the zero-mean projection matrix, $W=[\mathbf{w}_{1},\ldots,\mathbf{w}_{q}]\in\mathbb{R}^{d\times q}$, $V=[{\bm{v}}_{1},\ldots,{\bm{v}}_{q}]^{\top}\in\mathbb{R}^{d\times q}$ are learnable parameters. ${\bm{f}}[i]\in\mathbb{R}^{d}$ are input embeddings. $i$ is the sample index. Note that variants of quadratic activation have been used empirically, e.g. squared ReLU and gated activations (So et al., 2021; Shazeer, 2020; Zhang et al., 2024).
Input and Output. The input contains the two group elements $g_{1}[i]$ and $g_{2}[i]$, encoded as ${\bm{f}}[i]=U_{G_{1}}{\bm{e}}_{g_{1}[i]}+U_{G_{2}}{\bm{e}}_{g_{2}[i]}$, where $U_{G_{1}}$ and $U_{G_{2}}$ are column orthogonal embedding matrices. The output is the result $g_{1}[i]g_{2}[i]\in G$, encoded as the label $l[i]=g_{1}[i]g_{2}[i]$ to be predicted. We can extend our framework to group action prediction, in which $g_{2}$ may not be a group element but any object (e.g., a discrete state in reinforcement learning). See Appendix E for more details.
Let $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}\in\mathbb{C}^{d}$ be the scaled Fourier bases (or more formally, character function of the finite Abelian group $G$, see Appendix A). Then weight vector $\mathbf{w}_{j}$ and ${\bm{v}}_{j}$ can be written as:
$\mathbf{w}_{j}=U_{G_{1}}\sum_{k\neq 0}z_{akj}\bm{\phi}_{k}+U_{G_{2}}\sum_{k\neq 0}z_{bkj}\bm{\phi}_{k},\quad\quad{\bm{v}}_{j}=\sum_{k\neq 0}z_{ckj}\bar{\bm{\phi}}_{k}$ (2)
where ${\bm{z}}:=\{z_{pkj}\}$ are the complex coefficients, $p\in\{a,b,c\}$, $0\leq k<d$ and $j$ runs through $q$ hidden nodes. We exclude $\bm{\phi}_{0}\equiv 1$ because the constant bias term has been filtered out by the top-down gradient from the loss function. Since $\mathbf{w}_{j}$ and ${\bm{v}}_{j}$ are all real, the Hermitian constraints holds, i.e., $\overline{z_{ckj}}=\overline{\bm{\phi}^{*}_{k}{\bm{v}}_{j}}=\bm{\phi}^{*}_{-k}{\bm{v}}_{j}=z_{c,-k,j}$ (and similar for $z_{akj}$ and $z_{bkj}$). Leveraging the property of quadratic activation functions, we can write down the loss function analytically (see Appendix A):
@@
Note that for cyclic group $G$, the frequency $k$ is a mod-$d$ integer. For general Abelian group which can be decomposed into direct sum of cyclic groups according to Fundamental Theorem of Finite Abelian Groups (Diaconis, 1988), $k$ is a multidimensional frequency index. For convenience, we define $\phi_{-k}:=\overline{\phi}_{k}$ as the conjugate representation of $\phi_{k}$. Since weights $\mathbf{w}_{j}$ and ${\bm{v}}_{j}$ are all real, the Hermitian constraints holds, i.e., $\overline{z_{ckj}}=\overline{\bm{\phi}^{*}_{k}{\bm{v}}_{j}}=\bm{\phi}^{*}_{-k}{\bm{v}}_{j}=z_{c,-k,j}$ (and similar for $z_{akj}$ and $z_{bkj}$). Therefore, $z_{p,-k,j}=\bar{z}_{pkj}$, $r_{-k,-k,-k}=\bar{r}_{kkk}$ and $\ell$ is real and can be minimized.
@@ 4.2 The Monomial Potentials and its connection to semi-ring $\mathcal{Z}$
So what is the relationship between MPs, which are functions that map a weight ${\bm{z}}$ to a complex scalar, and the semi-ring structure of $\mathcal{Z}$? The following theorem tells that MPs are ring homomorphism, that is, these mappings respect addition and multiplication:
@@ 5.2 Composing Global Solutions
Remarks. To construct $\hat{\bm{z}}_{F6}$, in addition to ${\bm{z}}_{\mathrm{syn}}*{\bm{z}}_{\nu=1}$, we could use other pairs of single frequency solutions to achieve the same effects. For example, using ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$, where ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ is:
@@
where $|\alpha|=|\beta|=1$. ${\bm{z}}_{\mathrm{syn}}=\bm{\rho}({\bm{u}}_{\mathrm{syn}})$ is a special case of ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ when $\alpha=\beta=1$.
@@ Appendix A Decoupling $L_{2}$ Loss (Proof)
For convenience, we define $\phi_{-k}:=\overline{\phi}_{k}$ as the conjugate representation of $\phi_{k}$.
Let $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}\in\mathbb{C}^{d}$ be the vector that contains the value of the character function $\phi_{k}$. Then $\{\bm{\phi}_{k}\}$ form an orthogonal base in $\mathbb{C}^{d}$ and we can represent the weight vector $\mathbf{w}_{j}$ and ${\bm{v}}_{j}$ as the following:
$\mathbf{w}_{j}=U_{G_{1}}\sum_{k\neq 0}z_{akj}\bm{\phi}_{k}+U_{G_{2}}\sum_{k\neq 0}z_{bkj}\bm{\phi}_{k},\quad\quad{\bm{v}}_{j}=\sum_{k\neq 0}z_{ckj}\bar{\bm{\phi}}_{k}$ (12)
where ${\bm{z}}:=\{z_{pkj}\}$ are the complex coefficients ($p\in\{a,b,c\}$, $0\leq k<d$ and $j$ runs through hidden nodes). Then it is clear that $\mathbf{w}_{j}^{\top}{\bm{f}}[i]=\sum_{k\neq 0}h_{akj}\phi_{k}(\iota_{0}(g[i]))+\sum_{k\neq 0}h_{bkj}\phi_{k}(x[i])$.
@@
$\ell$ $=$ $\mathbb{E}_{g,x}\left[\|P_{1}^{\perp}({\bm{o}}(g,x)/2d-{\bm{e}}_{gx})\|^{2}\right]$ (13)
$=$ $\mathbb{E}_{g,x}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{o}}/4d^{2}-{\bm{o}}^{\top}P_{1}^{\perp}{\bm{e}}_{gx}/d+{\bm{e}}_{gx}^{\top}P_{1}^{\perp}{\bm{e}}_{gx}\right]$ (14)
For $\mathbb{E}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{e}}_{gx}\right]$, since
${\bm{e}}_{gx}^{\top}P_{1}^{\perp}{\bm{o}}$ $=$ $\sum_{j}{\bm{e}}_{gx}^{\top}P_{1}^{\perp}{\bm{v}}_{j}\sigma(\mathbf{w}_{j}^{\top}{\bm{f}}(g,x))$ (15)
$=$ $\sum_{j}\left(\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\phi}_{k^{\prime}}(gx)\right)\left(\sum_{k}a_{kj}\phi_{k}(\iota_{0}(g))+b_{kj}\phi_{k}(x)+{\bm{e}}^{\top}_{g}\mathbf{w}^{\perp}_{j}\right)^{2}$ (16)
Note that by our previous analysis, there exists $y_{1}:=\iota_{0}(g)$ so that $gy=x_{1}y$. Let $x_{2}:=x$. For notation brevity, let $z_{akj}:=a_{kj}$, $z_{bkj}:=b_{kj}$ and $z_{ckj}:=c_{kj}$,, then we have:
${\bm{e}}_{gx}^{\top}P_{1}^{\perp}{\bm{o}}=\sum_{j}\left(\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\phi}_{k^{\prime}}(x_{1}x_{2})\right)\left(\sum_{k}\sum_{p}z_{pkj}\phi_{k}(x_{p})+{\bm{e}}_{x_{1}}^{\top}\mathbf{w}_{j}^{\perp}\right)^{2}$ (17)
Therefore, we have:
$\mathbb{E}_{g,x}\left[{\bm{e}}_{gx}^{\top}P_{1}^{\perp}{\bm{o}}\right]=\sum_{k_{1},k_{2},k^{\prime}\neq 0,p_{1},p_{2},j}c_{k^{\prime}j}z_{p_{1}k_{1}j}z_{p_{2}k_{2}j}\mathbb{E}\left[\bar{\phi}_{k^{\prime}}(x_{1})\bar{\phi}_{k^{\prime}}(x_{2})\phi_{k_{1}}(x_{p_{1}})\phi_{k_{2}}(x_{p_{2}})\right]$ (18)
Note that due to the fact that $\mathbb{E}_{g\in\iota_{0}^{-1}(x_{1})}\left[{\bm{e}}_{g}^{\top}\mathbf{w}_{j}^{\perp}\right]=0$ and $\mathbb{E}_{g\in\iota_{0}^{-1}(x_{1})}\left[{\bm{e}}_{g}{\bm{e}}_{g}^{\top}\right]$ is only a function of $x_{1}$ and becomes 00 if multiplied with $\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\phi}_{k^{\prime}}(x_{1}x_{2})$ and taking expectation w.r.t $x_{2}$, in the final expression, all terms involving $\mathbf{w}_{j}^{\perp}$ vanish.
Since $\mathbb{E}_{x}\left[\phi_{k}(x)\bar{\phi}_{k^{\prime}}(x)\right]=\mathbb{I}(k=k^{\prime})$, there are only a few cases that the summand is nonzero:
  • $p_{1}=1$, $p_{2}=2$, $k^{\prime}=k_{1}=k_{2}\neq 0$.
  • $p_{1}=2$, $p_{2}=1$, $k^{\prime}=k_{1}=k_{2}\neq 0$.
In both cases, the summation reduces to $\sum_{k\neq 0,j}c_{kj}z_{1kj}z_{2kj}=\sum_{k\neq 0,j}c_{kj}a_{kj}b_{kj}$. Let $r_{k_{1}k_{2}k^{\prime}}:=\sum_{j}a_{k_{1}j}b_{k_{2}j}c_{k^{\prime}j}$, then we have
$\mathbb{E}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{e}}_{gy}\right]=2\sum_{k\neq 0,j}a_{kj}b_{kj}c_{kj}=2\sum_{k\neq 0}r_{kkk}$ (19)
For $\mathbb{E}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{o}}\right]$, if $\mathbf{w}^{\perp}_{j}=0$, then we have:
${\bm{o}}^{\top}P_{1}^{\perp}{\bm{o}}=\sum_{j,j^{\prime}}{\bm{v}}^{\top}_{j}P_{1}^{\perp}{\bm{v}}_{j^{\prime}}\sigma(\mathbf{w}_{j}^{\top}{\bm{f}}(g,y))\sigma(\mathbf{w}_{j^{\prime}}^{\top}{\bm{f}}(g,y))$ (20)
@@
${\bm{v}}_{j}^{\top}P_{1}^{\perp}{\bm{v}}_{j^{\prime}}=\left(\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\bm{\phi}}_{k^{\prime}}\right)^{\top}\left(\sum_{k^{\prime\prime}\neq 0}\bar{c}_{k^{\prime\prime}j^{\prime}}\bm{\phi}_{k^{\prime\prime}}\right)=d\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{c}_{k^{\prime}j^{\prime}}$ (21)
due to the fact that $\bar{\bm{\phi}}^{\top}_{k}\bm{\phi}_{k^{\prime}}=\sum_{y}\bar{\phi}_{k}(y)\phi_{k^{\prime}}(y)=d\mathbb{I}(k=k^{\prime})$.

### v3

@@ Abstract
We prove rich algebraic structures of the solution space for 2-layer neural networks with quadratic activation and $L_{2}$ loss, trained on reasoning tasks in Abelian group (e.g., modular addition). Such a rich structure enables analytical construction of global optimal solutions from partial solutions that only satisfy part of the loss, despite its high nonlinearity. We coin the framework as CoGO (Composing Global Optimizers). Specifically, we show that the weight space over different numbers of hidden nodes of the 2-layer network is equipped with a semi-ring algebraic structure, and the loss function to be optimized consists of monomial potentials, which are ring homomorphisms, allowing partial solutions to be composed into global ones by ring addition and multiplication. Our experiments show that around $95\%$ of the solutions obtained by gradient descent match exactly our theoretical constructions. Although the global optimizers constructed only required a small number of hidden nodes, our analysis on gradient dynamics shows that overparameterization asymptotically decouples training dynamics and is beneficial. We further show that training dynamics favors simpler solutions under weight decay, and thus high-order global optimizers such as perfect memorization are unfavorable. The code is at [^1] https://github.com/facebookresearch/luckmatters/tree/yuandong3/ssl/real-dataset.
@@ 1 Introduction
As a result, our theoretical framework, named CoGO (i.e., Composing Global Optimizers), successfully constructs two distinct types of Fourier-based global optimizers of per-frequency order 4 (or “$2\times 2$”) and order 6 (or “$2\times 3$”), and a global optimizer of order $d^{2}$ that correspond to perfect memorization. Empirically, we demonstrate that around $95\%$ of the solutions obtained from gradient descent (with weight decay) have the predicted structure and match exactly with our theoretical construction of order-4 and order-6 solutions. In addition, we also analyze the training dynamics, and show that the dynamics favors low-order global optimizers, since global optimizers algebraically connected by ring multiplication can be proven to also be topologically connected. Therefore, high-order solution like perfect memorization is unfavorable in the dynamics. When the network width goes to infinity, the dynamics of monomial potentials becomes decoupled, demystifying why overparameterization improves the performance.
@@ 2 Related Works
Figure 1: Overview of proposed theoretical framework CoGO. (1) The family of 2-layer neural networks, $\mathcal{Z}$, form a semi-ring algebraic structure (Theorem 2) with ring addition and multiplication (Def. 5). $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$ where $\mathcal{Z}_{q}$ is a collection of all weights with order-$q$ (i.e., $q$ hidden nodes). (2) For outcome prediction of Abelian group multiplication, the MSE loss $\ell({\bm{z}})$ is a function of monomial potentials (MPs) $r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$ (Theorem 1), which are ring homomorphisms (Theorem 3). (3) Thanks to the property of ring homomorphism, global optimizers to MSE loss $\ell({\bm{z}})$ with quadratic activation can be constructed algebraically from partial solutions that only satisfy a subset of constraints (Sec. 5.1) using ring addition and multiplication, instead of running gradient descent. Examples include Fourier solution ${\bm{z}}_{F6}$ (Corollary 2) and ${\bm{z}}_{F4/6}$ (Corollary 4) and perfect memorization solution ${\bm{z}}_{M}$ (Corollary 3). In Sec. 6, we analyze the role played of MPs in gradient dynamics, showing that the dynamics favors low-order global optimizers (Theorem 6) under weight decay regularization, and the dynamics of MPs become decoupled with infinite width (Theorem 7). In Sec. 7 we show that the gradient descent solutions match exactly with our theoretical construction.
@@ 3 Decoupling $L_{2}$ Loss for reasoning tasks of Abelian group
Basic Ring theory. A set $\mathcal{Z}$ forms a ring, if there exists two operations, addition $+$ and multiplication *, so that (1) $\langle\mathcal{Z},+\rangle$ forms an Abelian group, (2) $\langle\mathcal{Z},*\rangle$ is a monoid (i.e., a group without inverse), and (3) multiplication distributes with addition (i.e., $a*(b+c)=a*b+a*c$ and $(b+c)*a=b*a+c*a$). $\mathcal{Z}$ is called a semi-ring if $\langle\mathcal{Z},+\rangle$ is a monoid.
@@
$\ell=\sum_{i}\Big{\|}P^{\perp}_{1}\left(\frac{1}{2d}{\bm{o}}[i]-{\bm{e}}_{l[i]}\right)\Big{\|}^{2},\quad\quad{\bm{o}}[i]=\sum_{j}\mathbf{w}_{cj}\sigma(\mathbf{w}_{aj}^{\top}{\bm{e}}_{g_{1}[i]}+\mathbf{w}_{bj}^{\top}{\bm{e}}_{g_{2}[i]})$ (1)
Input and Output. The input contains the two group elements $g_{1}[i],g_{2}[i]\in G$ to be multiplied, ${\bm{e}}_{g_{1}[i]},{\bm{e}}_{g_{2}[i]}\in\mathbb{R}^{d}$ are one-hot representation of $g_{1}[i]$ and $g_{2}[i]$. Here $i$ is the sample index. The target ${\bm{e}}_{l[i]}$ is a one-hot representation of $l[i]=g_{1}[i]g_{2}[i]\in G$, the group product of $g_{1}[i]$ and $g_{2}[i]$.
Figure 2: Problem setup. For group multiplication $l[i]=g_{1}[i]g_{2}[i]\mod d$, we use a 2-layer network with quadratic activation to predict $l[i]$ from $g_{1}[i]$ and $g_{2}[i]$. See Eqn. 1 for the loss function.
Architectures. In Eqn. 1, we use quadratic activation $\sigma(x)=x^{2}$ (Du & Lee, 2018; Allen-Zhu & Li, 2023), $P_{1}^{\perp}=I-\frac{1}{d}{\bm{1}}{\bm{1}}^{\top}$ is the zero-mean projection, $\mathbf{w}_{aj},\mathbf{w}_{bj},\mathbf{w}_{cj}\in\mathbb{R}^{d}$ are learnable parameters ($1\leq j\leq q$). See Fig. 2. Note that variants of quadratic activation have been used empirically, e.g. squared ReLU and gated activations (So et al., 2021; Shazeer, 2020; Zhang et al., 2024).
We can extend our framework to group action prediction, in which $g_{2}$ may not be a group element but any object (e.g., a discrete state in reinforcement learning). See Appendix E for more details.
Let $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}\in\mathbb{C}^{d}$ be the scaled Fourier bases (or more formally, character function of the finite Abelian group $G$, see Appendix A). Then the weight vector $\mathcal{W}:=\{\mathbf{w}_{j}\}$ can be written as:
$\mathbf{w}_{aj}=\sum_{k\neq 0}z_{akj}\bm{\phi}_{k},\quad\quad\mathbf{w}_{bj}=\sum_{k\neq 0}z_{bkj}\bm{\phi}_{k},\quad\quad\mathbf{w}_{cj}=\sum_{k\neq 0}z_{ckj}\bar{\bm{\phi}}_{k}$ (2)
where ${\bm{z}}:=\{z_{pkj}\}$ are the complex coefficients, $p\in\{a,b,c\}$, $0\leq k<d$ and $j$ runs through $q$ hidden nodes. For convenience, we define $\bm{\phi}_{-k}:=\overline{\bm{\phi}}_{k}$ as the (complex) conjugate representation of $\bm{\phi}_{k}$. We exclude $\bm{\phi}_{0}\equiv 1$ because the constant bias term has been filtered out by the top-down gradient from the loss function. Leveraging the property of quadratic activation functions, we can write down the loss function analytically (see Appendix A):
@@
Note that for cyclic group $G$, the frequency $k$ is a mod-$d$ integer. For general Abelian group which can be decomposed into a direct sum of cyclic groups according to Fundamental Theorem of Finite Abelian Groups (Diaconis, 1988), $k$ is a multidimensional frequency index. Since $\{\mathbf{w}_{pj}\}$ are all real, the Hermitian constraints hold, i.e. $\overline{z_{ckj}}=\overline{\bm{\phi}^{*}_{k}\mathbf{w}_{cj}}=\bm{\phi}^{*}_{-k}\mathbf{w}_{cj}=z_{c,-k,j}$ (and similar for $z_{akj}$ and $z_{bkj}$). Therefore, $z_{p,-k,j}=\bar{z}_{pkj}$, $r_{-k,-k,-k}=\bar{r}_{kkk}$ and $\ell$ is real and can be minimized.
@@ 4.2 The Monomial Potentials and its connection to semi-ring $\mathcal{Z}$
So what is the relationship between MPs, which are functions that map a weight ${\bm{z}}$ to a complex scalar, and the semi-ring structure of $\mathcal{Z}$? The following theorem tells that MPs are ring homomorphisms, that is, these mappings respect addition and multiplication:
@@ 5.2 Composing Global Solutions
Remarks. To construct $\hat{\bm{z}}_{F6}$, in addition to ${\bm{z}}_{\mathrm{syn}}*{\bm{z}}_{\nu=1}$, we could use other compositions of single frequency solutions to achieve the same effects. For example, ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$, where ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ is:
@@
where $|\alpha|=|\beta|=1$. Note that ${\bm{z}}_{\mathrm{syn}}=\bm{\rho}({\bm{u}}_{\mathrm{syn}})$ is a special case of ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ when $\alpha=\beta=1$.
@@
Necessary Conditions. For now we have discussed solutions constructed to be global optimizers (i.e. sufficient conditions), For the structure that weights must follow to satisfy Eqn. 1 (i.e. necessary conditions), please check Lemma 5 and Lemma 6 in Appendix Sec. C.3.
@@ Appendix A Decoupling $L_{2}$ Loss (Proof)
For convenience, we define $\phi_{-k}:=\overline{\phi}_{k}$ as the (complex) conjugate representation of $\phi_{k}$.
Let $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}\in\mathbb{C}^{d}$ be the vector that contains the value of the character function $\phi_{k}$ over $G$. Then $\{\bm{\phi}_{k}\}$ form an orthogonal base in $\mathbb{C}^{d}$ and we can represent the weight vector $\mathbf{w}_{pj}$ as the following, where $p\in\{a,b,c\}$:
$\mathbf{w}_{aj}=\sum_{k\neq 0}z_{akj}\bm{\phi}_{k},\quad\quad\mathbf{w}_{bj}=\sum_{k\neq 0}z_{bkj}\bm{\phi}_{k},\quad\quad\mathbf{w}_{cj}=\sum_{k\neq 0}z_{ckj}\bar{\bm{\phi}}_{k}$ (12)
where ${\bm{z}}:=\{z_{pkj}\}$ are the complex coefficients. Here $p\in\{a,b,c\}$, $0\leq k<d$ and $j$ runs through hidden nodes.
@@
$\ell$ $=$ $\mathbb{E}_{g_{1},g_{2}}\left[\|P_{1}^{\perp}({\bm{o}}(g_{1},g_{2})/2d-{\bm{e}}_{g_{1}g_{2}})\|^{2}\right]$ (13)
$=$ $\mathbb{E}_{g_{1},g_{2}}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{o}}/4d^{2}-{\bm{o}}^{\top}P_{1}^{\perp}{\bm{e}}_{g_{1}g_{2}}/d+{\bm{e}}_{g_{1}g_{2}}^{\top}P_{1}^{\perp}{\bm{e}}_{g_{1}g_{2}}\right]$ (14)
For notation brevity, let $z_{akj}:=a_{kj}$, $z_{bkj}:=b_{kj}$ and $z_{ckj}:=c_{kj}$. For $\mathbb{E}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{e}}_{g_{1}g_{2}}\right]$, since
${\bm{e}}_{g_{1}g_{2}}^{\top}P_{1}^{\perp}{\bm{o}}$ $=$ $\sum_{j}{\bm{e}}_{g_{1}g_{2}}^{\top}P_{1}^{\perp}\mathbf{w}_{cj}\sigma(\mathbf{w}_{aj}^{\top}{\bm{e}}_{g_{1}}+\mathbf{w}_{bj}^{\top}{\bm{e}}_{g_{2}})$ (15)
$=$ $\sum_{j}\left(\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\phi}_{k^{\prime}}(g_{1}g_{2})\right)\left(\mathbf{w}_{aj}^{\top}{\bm{e}}_{g_{1}}+\mathbf{w}_{bj}^{\top}{\bm{e}}_{g_{2}}\right)^{2}$ (16)
$=$ $\sum_{j}\left(\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\phi}_{k^{\prime}}(g_{1}g_{2})\right)\left(\sum_{k}\sum_{p\in\{a,b\}}z_{pkj}\phi_{k}(g_{p})\right)^{2}$ (17)
Therefore, leveraging the fact that $\bar{\phi}_{k^{\prime}}(g_{1}g_{2})=\bar{\phi}_{k^{\prime}}(g_{1})\bar{\phi}_{k^{\prime}}(g_{2})$, we have:
$\mathbb{E}_{g_{1},g_{2}}\left[{\bm{e}}_{g_{1}g_{2}}^{\top}P_{1}^{\perp}{\bm{o}}\right]=\sum_{k_{1},k_{2},k^{\prime}\neq 0,p_{1},p_{2},j}c_{k^{\prime}j}z_{p_{1}k_{1}j}z_{p_{2}k_{2}j}\mathbb{E}_{g_{1},g_{2}}\left[\bar{\phi}_{k^{\prime}}(g_{1})\bar{\phi}_{k^{\prime}}(g_{2})\phi_{k_{1}}(g_{p_{1}})\phi_{k_{2}}(g_{p_{2}})\right]$ (18)
Since $\mathbb{E}_{g}\left[\phi_{k}(g)\bar{\phi}_{k^{\prime}}(g)\right]=\mathbb{I}(k=k^{\prime})$, there are only a few cases that the summand is nonzero:
  • $p_{1}=a$, $p_{2}=b$, $k^{\prime}=k_{1}=k_{2}\neq 0$.
  • $p_{1}=b$, $p_{2}=a$, $k^{\prime}=k_{1}=k_{2}\neq 0$.
In both cases, the summation reduces to $\sum_{k\neq 0,j}c_{kj}z_{akj}z_{bkj}=\sum_{k\neq 0,j}c_{kj}a_{kj}b_{kj}$. Let $r_{k_{1}k_{2}k^{\prime}}:=\sum_{j}a_{k_{1}j}b_{k_{2}j}c_{k^{\prime}j}$, then we have
$\mathbb{E}_{g_{1},g_{2}}\left[{\bm{o}}^{\top}(g_{1},g_{2})P_{1}^{\perp}{\bm{e}}_{g_{1}g_{2}}\right]=2\sum_{k\neq 0,j}a_{kj}b_{kj}c_{kj}=2\sum_{k\neq 0}r_{kkk}$ (19)
For $\mathbb{E}\left[{\bm{o}}^{\top}P_{1}^{\perp}{\bm{o}}\right]$, we have:
${\bm{o}}^{\top}P_{1}^{\perp}{\bm{o}}=\sum_{j,j^{\prime}}\mathbf{w}_{cj}^{\top}P_{1}^{\perp}\mathbf{w}_{cj^{\prime}}\sigma(\mathbf{w}_{aj}^{\top}{\bm{e}}_{g_{1}}+\mathbf{w}_{bj}^{\top}{\bm{e}}_{g_{2}})\sigma(\mathbf{w}_{aj^{\prime}}^{\top}{\bm{e}}_{g_{1}}+\mathbf{w}_{bj^{\prime}}^{\top}{\bm{e}}_{g_{2}})$ (20)
@@
$\mathbf{w}_{cj}^{\top}P_{1}^{\perp}\mathbf{w}_{cj^{\prime}}=\left(\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{\bm{\phi}}_{k^{\prime}}\right)^{\top}\left(\sum_{k^{\prime\prime}\neq 0}\bar{c}_{k^{\prime\prime}j^{\prime}}\bm{\phi}_{k^{\prime\prime}}\right)=d\sum_{k^{\prime}\neq 0}c_{k^{\prime}j}\bar{c}_{k^{\prime}j^{\prime}}$ (21)
due to the fact that $\bar{\bm{\phi}}^{\top}_{k}\bm{\phi}_{k^{\prime}}=\sum_{g}\bar{\phi}_{k}(g)\phi_{k^{\prime}}(g)=d\mathbb{I}(k=k^{\prime})$.